A string is stretched and fastened to two points at a distance 0

l

0 apart. Motion is started by

displacing the string in the form y = k sin πx

l

from which it is released at time t = 0. Show

that the displacement of any point on the string at a distance x from one end at time t is

given by k sin πx

l

cos πax

l


1
Expert's answer
2021-07-16T07:02:28-0400

Solution

The displacement y(x,t) of the string at a distance X from end 0 at time t is give by the following equation;

"\frac{d^2y}{dt^2}=a^2\frac{d^2y}{dx^2}" ..........(a)

We establish the boundary conditions using the given information.

Since the ends of the spring ,x=0 and x=l are fixed,they undergo zero displacement at any time t;

y(0,t)=0 for t"\geq" 0..........(1)

y(l,t)=0 for t"\geq" 0...........(2)

Since the spring is released from rest,at t =0 ,the velocity of every point on the string in the y-direction is zero,hence;

"\frac{dy}{dx}" (x,0)=0 for 0"\leq" x"\leq" l..............(3)

The initial displacement is in the form of y=ksin"\frac{πx}{l}" ,where y(x,0) is the initial displacement of any point x in the y-direction ,hence

y(x,0)=ksin"\frac{πx}{l}" for 0"\leq" x"\leq" l ..........(4)

(1),(2),(3) and (4) are the boundary conditions of the problem.

The general solutions of equation (a) with respect to the vibration of the string is

y(x,t)=(Acos px +Bsin px)(C cos pat +Dsin pat) ......(b)

A,B,C,D and p are arbitrary constants which can be found by applying the boundary conditions;

Using condition (1) in (b),we have

A(C cos pat + Dsin pat)=0 for all t"\geq"0

A=0

Using condition (2) in (b) ,we have

Since A=0

B sin pl(C cos pat +Dsin pat)=0,for all

t"\geq" 0;

Bsin pl=0

Either B or sin pl=0

If B = 0 ,y(x,t)=0 and the equation is meaningless,so

sin pl=0

pl=nπ

p="\frac{nπ}{l}"

Where n=0,1,2,3..."\infin"

Differentiate both sides of (b) with respect to t:

"\frac{dy}{dt}" (x,y)=(Bsin px)pa(-Csinpat+Dcos pat) ........(c)

Where p="\frac{nπ}{l}"

Using condition (3) on (c),we have

Bsin px.paD=0

But B"\ne" 0 and p"\ne" 0 ,we get D=0

With the found constants ,the equation (b) reduces to

y(x,t)=BC sin "\frac{nπx}{l}" Cos"\frac{nπat}{l}" where n=1,2,3....."\infin"

Taking BC=k,

The most general solution would be

y(x,t)="\displaystyle\sum_{n=1}^\infin" "\lambda_n"Sin"\frac{nπ}{l}"Cos"\frac{nπat}{l}" ....(d)

"\lambda_n" can be found using condition (4) in (d)

ksin"\frac{πx}{l}" ="\lambda_nsin\frac{nπx}{l}"

By comparison,

"\lambda_n" =k

"\frac{πx}{l}=\frac{nπx}{l}"

n=1

Using this values,the equation of displacement at any time on the string will be

y(x,t)=ksin"\frac{πx}{l}" Cos"\frac{πax}{l}"